OpenStudy (anonymous):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 3 over 4 power, all over 2 to the 1 over 2 power

OpenStudy (anonymous):

will medal and fan

Nnesha (nnesha):

\[ \huge\rm \frac{ 2^\frac{ 3 }{ 4 } }{ 2^\frac{ 1 }{ 2 }}\] exponent rule \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

Nnesha (nnesha):

ello 3 days ago post :3 and i'm still alive :o

Nnesha (nnesha):

 bases are same so you can move 1/2 exponent to the top when you do that sign would change \[\huge\rm 2^{\frac{ 3 }{ 4}-\frac{ 1 }{ 2 }}\]